OpenStudy (anonymous):

A box has 60 balls, of which 22 are red, 19 are blue, and 19 are yellow, as before. Balls are extracted at random but this time not returned to the box. What is the probability that after performing this test three times exactly two are blue and one is red?

OpenStudy (anonymous):

when i do the calculation i got the answer 1818/17110 but the book answer have 5643/17110, i think i did something wrong my calculation

OpenStudy (anonymous):

this what i did to get the calculation \[\frac{ \left(\begin{matrix}19 \\ 2\end{matrix}\right)* \left(\begin{matrix}22 \\ 1\end{matrix}\right)*\left(\begin{matrix}19 \\ 0\end{matrix}\right)}{ \left(\begin{matrix}60 \\ 3\end{matrix}\right)}\]

OpenStudy (amistre64):

what is the test being performed?

OpenStudy (anonymous):

combination

OpenStudy (amistre64):

What is the probability that after performing ***this test*** three times exactly two are blue and one is red? the questions feels like its missing information ... prolly from a different question?

OpenStudy (anonymous):

k A box has 50 balls, of which 16 are red, 19 are blue, and 15 are yellow. Balls are extracted at random then returned to the box. What is the probability that after performing this test three times exactly two are yellow and one is blue?

OpenStudy (anonymous):

similar but a bit different w/ a number

OpenStudy (amistre64):

experiment ... might be a better word

OpenStudy (anonymous):

i see

OpenStudy (amistre64):

A box has 60 balls, of which 22 are red, 19 are blue, and 19 are yellow, as before. Balls are extracted at random but this time not returned to the box. What is the probability that after performing this (process) three times exactly two are blue and one is red? rbb brb bbr 3*(22.19.18)/(60.59.58) maybe?

OpenStudy (anonymous):

doing that the number will be large

OpenStudy (amistre64):

1881/17110 is what that turns out to be

OpenStudy (amistre64):

P(bbr) or P(brb) or P(rbb) are the only ways we can pull 2bs and 1r in 3 pulls right?

OpenStudy (anonymous):

correct

OpenStudy (amistre64):

and after we pull them out the remaining amounts decrease by 1 22 for red, 19 and 18 for blues and the ball count reduces 60,59,58 for the product of the denominator ...

OpenStudy (anonymous):

is like this